# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ Abstract
It has been known that stochastic gravitational wave backgrounds (SGWBs) have anisotropies generated by squeezed-type tensor non-Gaussianities originating from scalar-tensor-tensor (STT) and tensor-tensor-tensor cubic interactions. While the squeezed tensor non-Gaussianities in the standard slow-roll inflation with the Bunch-Davies vacuum state are suppressed due to the so-called consistency relation, those in extended models with the violation of the consistency relation can be enhanced. Among such extended models, we consider the inflation model with the non-Bunch-Davies state that has been known to enhance the squeezed tensor non-Gaussianities. We explicitly formulate the primordial STT bispectrum induced during inflation in the context of Horndeski theory with the non-Bunch-Davies state and show that the induced SGWB anisotropies can be enhanced. We then discuss the detectability of those anisotropies in future gravitational wave experiments.
@@ 1 Introduction
Recently, several international collaborations for the Pulsar Timing Array (PTA) gravitational wave (GW) experiments such as North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [1, 2] have reported evidence of nHz stochastic gravitational wave background (SGWB) [3, 4, 5, 6]. Furthermore, future ground- and space-based missions at higher frequency bands, e.g., Cosmic Explorer (CE) [7], Einstein Telescope (ET) [8], Laser Interferometer Space Antenna (LISA) [9, 10], Taiji [11], the Taiji-LISA network [12, 13], Decihertz Gravitational Wave Observatory (DECIGO) [14, 15], the Big Bang Observer (BBO) [16], etc., are expected to report other evidence for the SGWB at different frequency ranges. While the most plausible source of the SGWB is the superposition of the GWs from astrophysical compact objects such as black holes (see, e.g., Ref. [17]), various cosmological origins of the SGWB, in particular, from the early universe with new physics have also been widely discussed (see, e.g., Ref. [18]). It is, therefore, worth discussing the possibilities of specifying the source of SGWB and testing new physics in the early universe with the PTA experiments and the ground- and space-based missions.
@@
As one of the things that can help identify the sources, the anisotropies of SGWB have been recently investigated [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. In addition to the anisotropies associated with the metric fluctuations in the late time universe (see, e.g., Refs. [19, 20, 24, 26]) known as the so-called Sachs-Wolfe (SW) and integral SW effects for the CMB anisotropies, [^1] See also Refs. [25, 31, 33, 34, 35, 36] where the anisotropies of secondary gravitational waves sourced by the late-time perturbation have been studied. the SGWB anisotropies can be sourced by the primordial scalar-tensor-tensor (STT) and tensor-tensor-tensor (TTT) cubic interactions in the squeezed configuration [21, 22, 23, 27, 28, 29, 30, 32]. In the present paper, we discuss the SGWB anisotropies induced by the primordial STT correlations generated during inflation. In the standard single scalar inflation, the consistency relation prohibits any enhancements of the primordial non-Gaussianities in the squeezed configuration [37]. On the other hand, in extended models with its violation, those non-Gaussianities can be enhanced, and then the SGWB anisotropies in such non-standard inflationary models have been studied in Refs. [21, 22, 23, 27, 28, 29, 30, 32]. Here, we consider inflation with non-Bunch-Davies states among the extended models with the consistency relation being violated.
@@
Inflation with the non-Bunch-Davies states has been shown to potentially enhance the primordial non-Gaussianities as a consequence of the violation of the consistency relation [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. In particular, the bispectra of the scalar and tensor perturbations are enhanced in proportion to powers of $k_{S}/k_{L}$ where $k_{S}$ and $k_{L}$ respectively denote the comoving wavenumber of the short- and long-wavelength modes in the squeezed limit, that is, $k_{L}/k_{S}\to 0$, and then it is naively expected that such a $k_{S}/k_{L}$ enhancement could efficiently amplify the SGWB anisotropies. However, since this characteristic enhancement of the primordial non-Gaussianities due to the non-Bunch-Davies states is basically induced by the cubic interactions on subhorizon scales, the short-wavelength modes may suffer from a so-called Trans-Planckian problem [79] in the squeezed limit $k_{S}\gg k_{L}$ when $k_{S}$ and $k_{L}$ are respectively taken to be the target frequency of observed SGWB and the scale of interest for their anisotropies on large scales. To obtain proper predictions for the SGWB anisotropies from the inflation with non-Bunch-Davies states, we first derive a condition to avoid the Trans-Planckian problem and then obtain the primordial STT bispectrum from the inflation with non-Bunch-Davies states under that condition. Also, towards testing the SGWB anisotropies from inflation, it is convenient to study those in a general framework of inflation, and hence we work with the Horndeski theory which is the most general scalar-tensor theory whose equations of motion are second order [80, 81, 82] (see also Ref. [83] for a review). As a result, we will show that the resultant non-Gaussianities and SGWB anisotropies from the general single-field inflation can be enhanced in a different way than the $k_{S}/k_{L}$ enhancement. We also discuss whether the SGWB anisotropies can be detected with actual gravitational wave experiments. The DECIGO/BBO experiments are designed to test inflation models predicting an almost scale-invariant tensor spectrum with the tensor-to-scalar ratio being $\mathcal{O}(10^{-2})$, while the recent NANOGrav data can be explained from inflation with a certain blue-tilted tensor spectrum. The power spectrum of the non-Bunch-Davies tensor modes generally has scale dependence via a Bogoliubov coefficient, and hence it is possible for inflation with the non-Bunch-Davies state to explain the recent results of nHz SGWB [84]. In light of these GW observations, we consider two configurations of the Bogoliubov coefficient. One is a scale-independent coefficient. This configuration leads to an almost scale-invariant tensor spectrum, the SGWB originating from which can potentially be tested by DECIGO/BBO. The other is a coefficient peaked at a certain scale $k=k_{*}$. The resultant tensor power spectrum has a peak at $k=k_{*}$ as well, and hence, properly choosing the value of $k_{*}$, the resultant GW can explain the recent PTA data. We would also like to investigate the cross-correlation between the cosmic microwave background (CMB) and the SGWB anisotropies [27].
@@
This paper is organized as follows. In the following section, we briefly review our setup and primordial power spectra of scalar and tensor perturbations with the non-Bunch-Davies states. In Sec. 3, we first explain the Trans-Planckian problem for the short-wavelength GWs and clarify the condition to avoid it. Then, we compute the primordial STT bispectrum with the non-Bunch-Davies states under that condition. In Sec. 2.2, we introduce two phenomenological models for the Bogoliubov coefficient such that the tensor power spectrum is nearly scale-invariant or has a peak at a certain scale. In Sec. 4, after reviewing the impacts of the primordial non-Gaussianities on the SGWB anisotropies, we evaluate the auto-correlation function of those and the cross-correlation one between those and CMB anisotropies. Then, we discuss the detectability of those correlation functions. The summary of this paper is drawn in Sec. 5.
@@ 2 Primordial power spectra with non-Bunch-Davies states
To explore the impact of primordial non-Gaussianities induced by the non-Bunch-Davies states on the SGWB, it is useful to formulate those in a general framework of single-field inflation. We work with the Horndeski theory [80, 81, 82], which is the most general scalar-tensor theory whose equations of motion are second order. The action takes the following form [82],
@@
$\quad+G_{5}(\phi,X)G_{\mu\nu}\nabla^{\mu}\nabla^{\nu}\phi-\frac{1}{6}G_{5X}[(\Box\phi)^{3}-3\Box\phi(\nabla_{\mu}\nabla_{\nu}\phi)^{2}+2(\nabla_{\mu}\nabla_{\nu}\phi)^{3}]\biggr{]},$ (2.1)
with $G_{i}(\phi,X)$ being arbitrary functions of the scalar field $\phi$ and its canonical kinetic term $X:=-g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi/2$. Imposing appropriate conditions on the arbitrary functions, the scalar field drives the slow-roll inflation as an inflaton [82].
The perturbed metric around a spatially flat Friedmann-Lemaître-Robertson-Walker spacetime under the unitary gauge in which the spatial fluctuation of the scalar field vanishes, that is, $\phi=\phi(t)$, is of the form
${\rm d}s^{2}=-N^{2}{\rm d}t^{2}+g_{ij}({\rm d}x^{i}+N^{i}{\rm d}t)({\rm d}x^{j}+N^{j}{\rm d}t),$ (2.2)
@@
$N$ $=1+\alpha,\ N_{i}=\partial_{i}\beta,$ (2.3)
$g_{ij}$ $=a^{2}e^{2\zeta}\biggl{(}\delta_{ij}+h_{ij}+\frac{1}{2}h_{ik}h_{kj}+\cdots\biggr{)},$ (2.4)
@@ 2.1 Quadratic action and primordial power spectra with non-Bunch-Davies states
Expanding the action Eq. (2.1) up to the quadratic order in the perturbations and substituting the solutions of $\alpha$ and $\beta$ derived from the constraint equations into the perturbed action, one obtains [82]
$S^{(2)}_{\zeta}$ $=\int{\rm d}t{\rm d}^{3}xa^{3}\biggl{[}\mathcal{G}_{S}\dot{\zeta}^{2}-\frac{\mathcal{F}_{S}}{a^{2}}(\partial_{i}\zeta)^{2}\biggr{]},$ (2.5)
$S^{(2)}_{h}$ $=\frac{1}{8}\int{\rm d}t{\rm d}^{3}xa^{3}\biggl{[}\mathcal{G}_{T}\dot{h}_{ij}^{2}-\frac{\mathcal{F}_{T}}{a^{2}}(\partial_{k}h_{ij})^{2}\biggr{]},$ (2.6)
where $\mathcal{G}_{S},~{}\mathcal{F}_{S},~{}\mathcal{G}_{T}$, and $\mathcal{F}_{T}$ are given in terms of $\phi$, $X$, the arbitrary functions $G_{i}(\phi,X)$ and their derivatives with respect to $\phi$ and $X$, and the explicit expressions of these functions are given in Appendix A. Throughout this paper, we assume a slow-roll inflation with $\mathcal{G}_{S},~{}\mathcal{F}_{S},~{}\mathcal{G}_{T},~{}\mathcal{F}_{T}\simeq{\rm const.}$ in a quasi-de Sitter spacetime ($H\simeq{\rm const.}$). Specifically, we introduce the slow-roll parameters as
$\epsilon:=-\frac{\dot{H}}{H^{2}},\ g_{S}:=\frac{\dot{\mathcal{G}}_{S}}{H\mathcal{G}_{S}},\ f_{S}:=\frac{\dot{\mathcal{F}}_{S}}{H\mathcal{F}_{S}},\ g_{T}:=\frac{\dot{\mathcal{G}}_{T}}{H\mathcal{G}_{T}},\ f_{T}:=\frac{\dot{\mathcal{F}}_{T}}{H\mathcal{F}_{T}},$ (2.7)
@@
$\zeta(t,{\bf x})$ $=\int\frac{{\rm d}^{3}k}{(2\pi)^{3}}\tilde{\zeta}(t,{\bf k})e^{i\bf k\cdot\bf x},$ (2.8)
$h_{ij}(t,{\bf x})$ $=\int\frac{{\rm d}^{3}k}{(2\pi)^{3}}\tilde{h}_{ij}(t,{\bf k})e^{i\bf k\cdot\bf x}.$ (2.9)
@@
$\tilde{\zeta}\to~{}\hat{\zeta}(t,{\bf k})$ $=u_{k}(t)\hat{a}_{\bf k}+u_{k}^{*}(t)\hat{a}^{\dagger}_{-{\bf k}},$ (2.10)
$\tilde{h}_{ij}\to~{}\hat{h}_{ij}(t,{\bf k})$ $=\sum_{s}\biggl{[}v_{k}(t)e^{(s)}_{ij}({\bf k})\hat{a}^{(s)}_{\bf k}+v^{*}_{k}(t)e^{(s)*}_{ij}(-{\bf k})\hat{a}^{(s)\dagger}_{-{\bf k}}\biggr{]},$ (2.11)
where $e^{(s)}_{ij}(\bf k)$ denotes the polarization tensor with the helicity states $s=\pm 2$, satisfying the transverse-traceless conditions, i.e., $k_{i}e^{(s)}_{ij}({\bf k})=\delta_{ij}e^{(s)}_{ij}({\bf k})=0$. The creation and annihilation operators satisfy the standard commutation relations,
$[\hat{a}_{\bf k},\hat{a}^{\dagger}_{{\bf k}^{\prime}}]$ $=(2\pi)^{3}\delta({\bf k}-{\bf k}^{\prime}),$ (2.12)
$[\hat{a}^{(s)}_{\bf k},\hat{a}^{(s^{\prime})\dagger}_{{\bf k}^{\prime}}]$ $=(2\pi)^{3}\delta_{ss^{\prime}}\delta({\bf k}-{\bf k}^{\prime}),$ (2.13)
${\rm others}$ $=0.$ (2.14)
@@
$E^{s}$ $:=\partial_{t}(a^{3}\mathcal{G}_{S}\dot{\zeta})-a\mathcal{F}_{S}\partial^{2}\zeta=0,$ (2.15)
$E^{h}_{ij}$ $:=\partial_{t}(a^{3}\mathcal{G}_{T}\dot{h}_{ij})-a\mathcal{F}_{T}\partial^{2}h_{ij}=0.$ (2.16)
@@
$u_{k}$ $=\alpha_{k}u_{k,({\rm BD})}+\beta_{k}u^{*}_{k,({\rm BD})},$ (2.17)
$v^{(s)}_{k}$ $=\alpha^{(s)}_{k}v_{k,({\rm BD})}+\beta^{(s)}_{k}v^{*}_{k,({\rm BD})},$ (2.18)
where $\alpha_{k},~{}\beta_{k},~{}\alpha^{(s)}_{k}$, and $\beta^{(s)}_{k}$ are Bogoliubov coefficients, and $u_{k,({\rm BD})}$ and $v_{k,({\rm BD})}$ denote the mode functions with the Bunch-Davies state which are respectively given as [82]
$u_{k,({\rm BD})}$ $=\frac{iH}{2\sqrt{\mathcal{F}_{S}c_{s}k^{3}}}(1+ic_{s}k\eta)e^{-ic_{s}k\eta},$ (2.19)
$v_{k,({\rm BD})}$ $=\frac{i\sqrt{2}H}{\sqrt{\mathcal{F}_{T}c_{h}k^{3}}}(1+ic_{h}k\eta)e^{-ic_{h}k\eta},$ (2.20)
@@
$|\alpha_{k}|^{2}-|\beta_{k}|^{2}$ $=1,$ (2.21)
$|\alpha^{(s)}_{k}|^{2}-|\beta^{(s)}_{k}|^{2}$ $=1,$ (2.22)

### v3

@@ Abstract
It is known that stochastic gravitational wave backgrounds (SGWBs) have anisotropies generated by squeezed-type tensor non-Gaussianities originating from scalar-tensor-tensor (STT) and tensor-tensor-tensor cubic interactions. While the squeezed tensor non-Gaussianities in the standard slow-roll inflation with the Bunch-Davies vacuum state are suppressed due to the so-called consistency relation, those in extended models with the violation of the consistency relation can be enhanced. Among such extended models, we consider the inflation model with the non-Bunch-Davies state that is known to enhance the squeezed tensor non-Gaussianities. We explicitly formulate the primordial STT bispectrum induced during inflation in the context of Horndeski theory with the non-Bunch-Davies state and show that the induced SGWB anisotropies can be enhanced. We then discuss the detectability of those anisotropies in future gravitational wave experiments.
[…]
E02,E03,E81
@@ 1 Introduction
Recently, several international collaborations for the Pulsar Timing Array (PTA) gravitational wave (GW) experiments such as North American Nanohertz Observatory for Gravitational Waves (NANOGrav) McLaughlin:2013ira ; Brazier:2019mmu have reported evidence of nHz stochastic gravitational wave background (SGWB) NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Xu:2023wog . Furthermore, future ground- and space-based missions at higher frequency bands, e.g. Cosmic Explorer (CE) Reitze:2019iox , Einstein Telescope (ET) Punturo:2010zz , Laser Interferometer Space Antenna (LISA) LISA:2017pwj ; LISACosmologyWorkingGroup:2022jok , Taiji Hu:2017mde , the Taiji-LISA network Ruan:2020smc ; Wang:2021uih , Decihertz Gravitational Wave Observatory (DECIGO) Kawamura:2006up ; Kawamura:2020pcg , the Big Bang Observer (BBO) Crowder:2005nr , etc., are expected to report other evidence for the SGWB at different frequency ranges. Although the most plausible source of the SGWB is the superposition of the GWs from astrophysical compact objects such as black holes (see, e.g. Ref. NANOGrav:2023hfp ), various cosmological origins of the SGWB, in particular, from the early universe with new physics have also been widely discussed (see, e.g. Ref. NANOGrav:2023hvm ). It is, therefore, worth discussing the possibilities of specifying the source of SGWB and testing new physics in the early universe with the PTA experiments and the ground- and space-based missions.
@@
As one of the things that can help identify the sources, the anisotropies of SGWB have been recently investigated Contaldi:2016koz ; Cusin:2017fwz ; Ricciardone:2017kre ; Dimastrogiovanni:2018uqy ; Dimastrogiovanni:2019bfl ; Bartolo:2019oiq ; Bartolo:2019zvb ; Bartolo:2019yeu ; Adshead:2020bji ; Malhotra:2020ket ; Dimastrogiovanni:2021mfs ; Dimastrogiovanni:2022afr ; Dimastrogiovanni:2022eir ; Orlando:2022rih ; Malhotra:2022ply ; Li:2023qua ; Li:2023xtl ; Yu:2023jrs . In addition to the anisotropies associated with the metric fluctuations in the late time universe (see, e.g. Refs. Contaldi:2016koz ; Cusin:2017fwz ; Bartolo:2019oiq ; Bartolo:2019yeu ) known as the so-called Sachs-Wolfe (SW) and integral SW effects for the cosmic microwave background (CMB) anisotropies, [^1] See also Refs. Bartolo:2019zvb ; Dimastrogiovanni:2022eir ; Malhotra:2022ply ; Li:2023qua ; Li:2023xtl ; Yu:2023jrs where the anisotropies of secondary GWs sourced by the late-time perturbation have been studied. the SGWB anisotropies can be sourced by the primordial scalar-tensor-tensor (STT) and tensor-tensor-tensor (TTT) cubic interactions in the squeezed configuration Ricciardone:2017kre ; Dimastrogiovanni:2018uqy ; Dimastrogiovanni:2019bfl ; Adshead:2020bji ; Malhotra:2020ket ; Dimastrogiovanni:2021mfs ; Dimastrogiovanni:2022afr ; Orlando:2022rih . In the present paper, we discuss the SGWB anisotropies induced by the primordial STT correlations generated during inflation. In the standard single scalar inflation, the consistency relation prohibits any enhancements of the primordial non-Gaussianities in the squeezed configuration Maldacena:2002vr . On the other hand, in extended models with its violation, those non-Gaussianities can be enhanced, and then the SGWB anisotropies in such non-standard inflationary models have been studied in Refs. Ricciardone:2017kre ; Dimastrogiovanni:2018uqy ; Dimastrogiovanni:2019bfl ; Adshead:2020bji ; Malhotra:2020ket ; Dimastrogiovanni:2021mfs ; Dimastrogiovanni:2022afr ; Orlando:2022rih . Here, we consider inflation with non-Bunch-Davies states among the extended models wherein the consistency relation is violated.
@@
Inflation with the non-Bunch-Davies states has been shown to potentially enhance the primordial non-Gaussianities as a consequence of the violation of the consistency relation Chen:2006nt ; Holman:2007na ; Xue:2008mk ; Meerburg:2009ys ; Meerburg:2009fi ; Chen:2010xka ; Meerburg:2010ks ; Chen:2010bka ; Meerburg:2010rp ; Agullo:2010ws ; Ashoorioon:2010xg ; Ganc:2011dy ; LopezNacir:2011kk ; Ganc:2012ae ; Agarwal:2012mq ; Gong:2013yvl ; Flauger:2013hra ; Aravind:2013lra ; Ashoorioon:2013eia ; Brahma:2013rua ; Bahrami:2013isa ; Kundu:2013gha ; Emami:2014tpa ; Zeynizadeh:2015eia ; Meerburg:2015yka ; Ashoorioon:2016lrg ; Shukla:2016bnu ; Ashoorioon:2018sqb ; Brahma:2019unn ; Akama:2020jko ; Naskar:2020vkd ; Ragavendra:2020vud ; Kanno:2022mkx ; Ghosh:2022cny ; Gong:2023kpe ; Ghosh:2023agt ; Akama:2023jsb ; Cielo:2023enz ; Peng:2024eok ; Ansari:2024pgq ; Christodoulidis:2024ric . In particular, the bispectra of the scalar and tensor perturbations are enhanced in proportion to powers of $k_{S}/k_{L}$ where $k_{S}$ and $k_{L}$ respectively denote the comoving wavenumber of the short- and long-wavelength modes in the squeezed limit, i.e. $k_{L}/k_{S}\to 0$, and then it is naively expected that such a $k_{S}/k_{L}$ enhancement could efficiently amplify the SGWB anisotropies. However, since this characteristic enhancement of the primordial non-Gaussianities due to the non-Bunch-Davies states is basically induced by the cubic interactions on subhorizon scales, the short-wavelength modes may suffer from a so-called Trans-Planckian problem Martin:2000xs and a strong coupling problem in the squeezed limit $k_{S}\gg k_{L}$ when $k_{S}$ and $k_{L}$ are respectively taken to be the target frequency of observed SGWBs and the scale of interest for their anisotropies on large scales. To obtain proper predictions for the SGWB anisotropies from the inflation with non-Bunch-Davies states, we first derive a condition to avoid the Trans-Planckian and strong coupling problems and then obtain the primordial STT bispectrum from the inflation with non-Bunch-Davies states under that condition. Also, for testing the SGWB anisotropies from inflation, it is convenient to study those in a general framework of inflation, and hence we work with the Horndeski theory, which is the most general scalar-tensor theory whose equations of motion are second order Horndeski:1974wa ; Deffayet:2011gz ; Kobayashi:2011nu (see also Ref. Kobayashi:2019hrl for a review). As a result, we will show that the resultant non-Gaussianities and SGWB anisotropies from the general single-field inflation can be enhanced in a different way than the $k_{S}/k_{L}$ enhancement. We also discuss whether the SGWB anisotropies can be detected with actual gravitational wave experiments. The DECIGO/BBO experiments are designed to test inflation models predicting an almost scale-invariant tensor spectrum with the tensor-to-scalar ratio being $\mathcal{O}(10^{-2})$, whereas the recent NANOGrav data can be explained from inflation with a certain blue-tilted tensor spectrum. The power spectrum of the non-Bunch-Davies tensor modes generally has scale dependence via a Bogoliubov coefficient, and hence it is possible for inflation with the non-Bunch-Davies states to explain the recent results of nHz SGWB Choudhury:2023kam . In light of these GW observations, we consider two configurations of the Bogoliubov coefficient. One is a scale-independent coefficient. This configuration leads to an almost scale-invariant tensor spectrum; the SGWB originating from which can potentially be tested by DECIGO/BBO. The other is a coefficient that peaks at a certain scale $k=k_{*}$. The resultant tensor power spectrum has a peak at $k=k_{*}$ as well, and hence, properly choosing the value of $k_{*}$, the resultant GW can explain the recent PTA data. We would also like to investigate the cross-correlation between the CMB and the SGWB anisotropies Adshead:2020bji .
@@
This paper is organized as follows. In the following section, we briefly review our setup and primordial power spectra of scalar and tensor perturbations with the non-Bunch-Davies states. In Sec. 3, we first explain the Trans-Planckian and strong coupling problems for the short-wavelength GWs and clarify the condition to avoid these problems. Then, we compute the primordial STT bispectrum with the non-Bunch-Davies states under that condition. In Sec. 2.2, we introduce two phenomenological models for the Bogoliubov coefficient such that the tensor power spectrum is nearly scale-invariant or has a peak at a certain scale. In Sec. 4, after reviewing the impacts of the primordial non-Gaussianities on the SGWB anisotropies, we evaluate the auto-correlation function of those and the cross-correlation one between those and CMB anisotropies. Then, we discuss the detectability of those correlation functions. The summary of this paper is drawn in Sec. 5.
@@ 2 Primordial power spectra with non-Bunch-Davies states
To explore the impact of primordial non-Gaussianities induced by the non-Bunch-Davies states on the SGWB, it is useful to formulate them in a general framework of single-field inflation. We work with the Horndeski theory Horndeski:1974wa ; Deffayet:2011gz ; Kobayashi:2011nu , which is the most general scalar-tensor theory whose equations of motion are second order. The action takes the following form Kobayashi:2011nu :
@@
$\quad+G_{5}(\phi,X)G_{\mu\nu}\nabla^{\mu}\nabla^{\nu}\phi-\frac{1}{6}G_{5X}[(\Box\phi)^{3}-3\Box\phi(\nabla_{\mu}\nabla_{\nu}\phi)^{2}+2(\nabla_{\mu}\nabla_{\nu}\phi)^{3}]\biggr],$ (1)
with $G_{i}(\phi,X)$ being arbitrary functions of the scalar field $\phi$ and its canonical kinetic term $X:=-g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi/2$. Imposing appropriate conditions on the arbitrary functions, the scalar field drives the slow-roll inflation as an inflaton Kobayashi:2011nu .
The perturbed metric around a spatially flat Friedmann-Lemaître-Robertson-Walker spacetime under the unitary gauge in which the spatial fluctuation of the scalar field vanishes, i.e. $\phi=\phi(t)$, is of the form
${\rm d}s^{2}=-N^{2}{\rm d}t^{2}+g_{ij}({\rm d}x^{i}+N^{i}{\rm d}t)({\rm d}x^{j}+N^{j}{\rm d}t),$ (2)
@@
$N$ $=1+\alpha,\ N_{i}=\partial_{i}\beta,$ (3)
$g_{ij}$ $=a^{2}e^{2\zeta}\biggl(\delta_{ij}+h_{ij}+\frac{1}{2}h_{ik}h_{kj}+\cdots\biggr),$ (4)
@@ 2.1 Quadratic action and primordial power spectra with non-Bunch-Davies states
Expanding the action Eq. (1) up to the quadratic order in the perturbations and substituting the solutions of $\alpha$ and $\beta$ derived from the constraint equations into the perturbed action, one obtains Kobayashi:2011nu :
$S^{(2)}_{\zeta}$ $=\int{\rm d}t{\rm d}^{3}xa^{3}\biggl[\mathcal{G}_{S}\dot{\zeta}^{2}-\frac{\mathcal{F}_{S}}{a^{2}}(\partial_{i}\zeta)^{2}\biggr],$ (5)
$S^{(2)}_{h}$ $=\frac{1}{8}\int{\rm d}t{\rm d}^{3}xa^{3}\biggl[\mathcal{G}_{T}\dot{h}_{ij}^{2}-\frac{\mathcal{F}_{T}}{a^{2}}(\partial_{k}h_{ij})^{2}\biggr],$ (6)
where $\mathcal{G}_{S},~\mathcal{F}_{S},~\mathcal{G}_{T}$, and $\mathcal{F}_{T}$ are given in terms of $\phi$, $X$, the arbitrary functions $G_{i}(\phi,X)$, and their derivatives with respect to $\phi$ and $X$, and the explicit expressions of these functions are given in Appendix A. Throughout this paper, we assume a slow-roll inflation with $\mathcal{G}_{S},~\mathcal{F}_{S},~\mathcal{G}_{T},~\mathcal{F}_{T}\simeq{\rm const.}$ in a quasi-de Sitter spacetime ($H\simeq{\rm const.}$). Specifically, we introduce the slow-roll parameters as
$\epsilon:=-\frac{\dot{H}}{H^{2}},\ g_{S}:=\frac{\dot{\mathcal{G}}_{S}}{H\mathcal{G}_{S}},\ f_{S}:=\frac{\dot{\mathcal{F}}_{S}}{H\mathcal{F}_{S}},\ g_{T}:=\frac{\dot{\mathcal{G}}_{T}}{H\mathcal{G}_{T}},\ f_{T}:=\frac{\dot{\mathcal{F}}_{T}}{H\mathcal{F}_{T}},$ (7)
@@
$\zeta(t,{\bf x})$ $=\int\frac{{\rm d}^{3}k}{(2\pi)^{3}}\tilde{\zeta}(t,{\bf k})e^{i\bf k\cdot\bf x},$ (8)
$h_{ij}(t,{\bf x})$ $=\int\frac{{\rm d}^{3}k}{(2\pi)^{3}}\tilde{h}_{ij}(t,{\bf k})e^{i\bf k\cdot\bf x}.$ (9)
@@
$\tilde{\zeta}\to~\hat{\zeta}(t,{\bf k})$ $=u_{k}(t)\hat{a}_{\bf k}+u_{k}^{*}(t)\hat{a}^{\dagger}_{-{\bf k}},$ (10)
$\tilde{h}_{ij}\to~\hat{h}_{ij}(t,{\bf k})$ $=\sum_{s}\biggl[v_{k}(t)e^{(s)}_{ij}({\bf k})\hat{a}^{(s)}_{\bf k}+v^{*}_{k}(t)e^{(s)*}_{ij}(-{\bf k})\hat{a}^{(s)\dagger}_{-{\bf k}}\biggr$], (11)
where $e^{(s)}_{ij}(\bf k)$ denotes the polarization tensor with the helicity states $s=\pm 2$, satisfying the transverse-traceless conditions, i.e. $k_{i}e^{(s)}_{ij}({\bf k})=\delta_{ij}e^{(s)}_{ij}({\bf k})=0$. The creation and annihilation operators satisfy the standard commutation relations,
$[\hat{a}_{\bf k},\hat{a}^{\dagger}_{{\bf k}^{\prime}}]$ $=(2\pi)^{3}\delta({\bf k}-{\bf k}^{\prime}),$ (12)
$[\hat{a}^{(s)}_{\bf k},\hat{a}^{(s^{\prime})\dagger}_{{\bf k}^{\prime}}$] $=(2\pi)^{3}\delta_{ss^{\prime}}\delta({\bf k}-{\bf k}^{\prime}),$ (13)
${\rm others}$ $=0.$ (14)
@@
$E^{s}$ $:=\partial_{t}(a^{3}\mathcal{G}_{S}\dot{\zeta})-a\mathcal{F}_{S}\partial^{2}\zeta=0,$ (15)
$E^{h}_{ij}$ $:=\partial_{t}(a^{3}\mathcal{G}_{T}\dot{h}_{ij})-a\mathcal{F}_{T}\partial^{2}h_{ij}=0.$ (16)
@@
$u_{k}$ $=\alpha_{k}u_{k,({\rm BD})}+\beta_{k}u^{*}_{k,({\rm BD})},$ (17)
$v^{(s)}_{k}$ $=\alpha^{(s)}_{k}v_{k,({\rm BD})}+\beta^{(s)}_{k}v^{*}_{k,({\rm BD})},$ (18)
where $\alpha_{k},~\beta_{k},~\alpha^{(s)}_{k}$, and $\beta^{(s)}_{k}$ are Bogoliubov coefficients, and $u_{k,({\rm BD})}$ and $v_{k,({\rm BD})}$ denote the mode functions with the Bunch-Davies state which are respectively given as Kobayashi:2011nu :
$u_{k,({\rm BD})}$ $=\frac{iH}{2\sqrt{\mathcal{F}_{S}c_{s}k^{3}}}(1+ic_{s}k\eta)e^{-ic_{s}k\eta},$ (19)
$v_{k,({\rm BD})}$ $=\frac{i\sqrt{2}H}{\sqrt{\mathcal{F}_{T}c_{h}k^{3}}}(1+ic_{h}k\eta)e^{-ic_{h}k\eta},$ (20)
@@
$|\alpha_{k}|^{2}-|\beta_{k}|^{2}$ $=1,$ (21)
$|\alpha^{(s)}_{k}|^{2}-|\beta^{(s)}_{k}|^{2}$ $=1,$ (22)